OpenStudy (anonymous):

Write a direct variation equation that relates x to y. Then solve. If y = 8 when x = 4, find y when x = 5.

OpenStudy (unklerhaukus):

\[y(x)\propto x\]\[y(x)=kx\]\[y(4)=4k=8\]\[k=?\]\[y(5)=5\times?\]

OpenStudy (anonymous):

y∝ x so y1/x1=y2/x2 8/4=y2/5 y2=10